OpenStudy (anonymous):

Write three equivalent ratios for each ratio. 19 to 38.

OpenStudy (anonymous):

1 to 2

OpenStudy (anonymous):

19/38 = 1/2 = 2/4 = 3/6

OpenStudy (anonymous):

o wait im doin the wrong one

OpenStudy (anonymous):

its a different problem lol

OpenStudy (anonymous):

lol